(5 pts) Find the area of the region bounded by the graphs of y + 2 and y = [ +1,0 < x < 2. 2 Sketch the region.

show all work please
Answer #1

Solution y = x + 2 y = x+1 o<ne 2 y y = kite y y=n+1 2 -2 Area = . - S (y2-yi) du a=o, b=2, Y = 23 + 2, Y,= ht! +)-(14) 4, =